OpenStudy (anonymous):

COULD SOME PLEASE HELP AND Explain how you would find the product of 20^12 and 20^8. Include either the exponent rule or the expanded method in your explanation. Give your answer in exponential form.

OpenStudy (anonymous):

Multiplying powers of the same base : add the exponents.

OpenStudy (phi):

remember that 20^12 means 20*20*20*... *20 (20 times itself 12 times) and 20^8 means multiply 20 times itself 8 times so just write 20 times itself 12 times, and then times 20 times itself 8 times now count the number of 20's

OpenStudy (amistre64):

20 = 2(10) 20^(12) = (2*10)^12 = 2^12 * 10^12 might be useful